a particular planet has a radius of 1000km and the acceleration due to gravity at its surface is 4m/s^2. what is the acceleration due to gravity 1000km above its surface?
Answer #1

Here,

R = 1000 km

as go = 4 m/s^2

for the acceleration due to gravity

g = go/(1 + h/R)^2

g = 4/(1 + 1000/1000)^2

g = 4/4 = 1 m/s^2

the acceleration due to gravity is 1 m/s^2